OpenStudy (anonymous):

Write the equation of the line that passes through the given point and is parallel or perpendicular to the given line. Write the answer in slope-intercept form. (−4, 2), parallel to y + 7x = −20 please explain how to work this

OpenStudy (abb0t):

Rearrange it into standard form: y = mx+b m is your slope b is your y-coordinate value

OpenStudy (anonymous):

so how do set it up? where does the -4 and 2 go?

OpenStudy (abb0t):

plug it in for x, and y...

OpenStudy (abb0t):

You could also use Calculus and find the derivative of the function.

OpenStudy (anonymous):

i need to know step by step how to work this i dont understand

terenzreignz (terenzreignz):

Firstly, what's the slope of y + 7x = -20 ?

OpenStudy (anonymous):

7 ?

terenzreignz (terenzreignz):

No. First, arrange it in this form y = mx + b It's important that y is alone on one side, with no coefficients or negative signs. Once it is in this arrangement, whatever is the coefficient of x, that's the slope. So again, what's the slope of y + 7x = -20 ?

OpenStudy (anonymous):

is it -7 ?